# Vosh: Voxel-Mesh Hybrid Representation for Real-Time View Synthesis

### v1

@@ 6.2. Rendering speed comparison
We evaluated the performance of the methods applicable for rendering on mobile devices, including SNeRG++ (Hedman et al., 2021), MERF (Reiser et al., 2023), BakedSDF (Yariv et al., 2023), MobileNeRF (Chen et al., 2023), and NeRF2Mesh (Tang et al., 2023). For SNeRG++ and MERF, we follow their configurations with progressive upsampling turned off. We conduct testing on the iPhone 14 Pro equipped with the A16 chip with a resolution of 344*707 and on the Legion Y7000P 2019, which features NVIDIA GeForce GTX 1660Ti (Mobile version, 80W) and operates with a resolution of 1920*1080.
@@
Rendering speed tests for all comparative methods were conducted on the Mip-Nerf 360 dataset (Barron et al., 2022) in three outdoor scenes, using identical device settings, resolutions, and viewpoints. As shown on the right side of Table 1, our method can run with the real-time frame rate (¿30 FPS) on both iPhone 14 Pro and GTX 1660Ti. Particularly on the 1660Ti, our method demonstrates a considerable speed improvement compared to MERF and SNeRG. The latter two methods experience slow rendering speeds due to the necessity to sample each alive voxel traversed by every ray. BakedSDF is difficult to render on mobile phones and laptops with limited VRAM due to its substantial assets. MobileNeRF and NeRF2Mesh achieve rendering speeds that reach the maximum refresh rates of their respective devices (60Hz for Safari on iPhone 14 Pro and 144Hz for Legion Y7000P 2019). We make statistics on their potential rendering speeds using overlapping multiple passes.
6.3. Ablation Study
@@
We conduct ablation experiments on the voxels to mesh conversion and voxel adjustment parts of the hybrid representation in Table 2. The assessment involves a quantitative comparison, considering both rendering quality and speed.
The ‘full bound’ row refers to the results of the variant that do not use coarse selection of voxels in the voxels to mesh conversion. The experimental results show that converting all voxels containing surfaces to mesh not only reduces the rendering quality of the final hybrid representation, but also significantly increases the storage of the assets.
@@
In the voxel adjustment stage, we compare the complete model with variants of Vosh-Base, incorporating different loss weights or resolutions for mesh occupancy grids. The rows with $\lambda_{voxel}$ show the results of different hyper-parameter weights of the loss function for voxel adjustment. The rows with $r_{mesh}$ indicate the results of using mesh occupancy grids with different resolutions. For the Vosh-Base model, we use $\lambda=0.001$ and $r_{mesh}=128$ as hyper-parameters. For the Vosh-Light model, we use $\lambda=0.1$ and $r_{mesh}=32$ as hyper-parameters. Experimental results suggest that increasing the weight of voxel adjustment loss or reducing the resolution of the mesh occupancy grid aids in compressing the number of voxels, leading to speed benefits at the expense of rendering quality.
Table 2. The ablation study in the construction of hybrid representation.
Variants PSNR ↑ SSIM ↑ LPIPS ↓ FPS ↑
Vosh-Base 23.74 0.592 0.389 33.7
full bound 23.49 0.583 0.395 36.0
$\lambda_{voxel}=0$ 23.73 0.591 0.389 30.0
$\lambda_{voxel}=0.01$ 23.72 0.590 0.389 38.7
$\lambda_{voxel}=0.1$ 23.69 0.590 0.390 46.7
$r_{mesh}=0$ 23.80 0.593 0.386 32.0
$r_{mesh}=1024$ 23.79 0.593 0.387 32.0
$r_{mesh}=64$ 23.70 0.590 0.389 37.7
Vosh-Light 23.63 0.586 0.392 70.0
6.4. Limitations

### v2

@@ VI-C Rendering speed comparison
We evaluated the performance of the methods applicable for rendering on platform-independent mobile devices, including SNeRG++ [8], MERF [18], BakedSDF [16], MobileNeRF [13], and NeRF2Mesh [14]. We conduct testing on iPhone 14 Pro equipped with the A16 chip with a resolution of 347$\times$600 and on Lenovo Yoga 14s 2021, which features Intel Iris Xe Graphics 80EU (28W, integrated in Intel Core i5-11350H) and operates with a resolution of 800$\times$600. For SNeRG++ and MERF, we follow their configurations with progressive upsampling turned off.
@@
Rendering speed tests for all comparative methods are conducted on the Mip-Nerf 360 dataset [38] in three outdoor scenes, using identical device settings, resolutions, and viewpoints. As shown on the right side of Table I, our method can run with the real-time frame rate (>30 FPS) on both iPhone 14 Pro and Iris Xe 80EU. Adaptive Shell [19] can only achieve real-time rendering at 1080p resolution using an RTX 4090 GPU (36.18 FPS from its paper). Our method can achieve faster frame rates at the same resolution using an GTX 1660 Ti GPU. The volume-base methods have slow rendering speeds due to the necessity to sample each alive voxel traversed by every ray. BakedSDF [16] is difficult to render on mobile phones and laptops with limited VRAM due to its substantial assets. Specifically, BakedSDF has about 10 million triangles to represent unbounded scenes, mostly for complex backgrounds. Vosh, leveraging the efficiency of voxel component, reduces the need for triangles in low-quality and space-consuming background areas, using only about one-tenth as many. This greatly shortens the time of rasterization, providing an opportunity to enhance the performance based on sparse voxels. NeRF2Mesh [14] achieves rendering speeds that reach the maximum refresh rates of their respective devices (60Hz for Safari on iPhone 14 Pro and 90Hz for Lenovo Yoga 14s 2021). We make statistics on their potential rendering speeds using overlapping multiple passes.
VI-D Comparison with VMesh
It is worth noting that VMesh [20] is also built on the hybrid representation of volume and mesh, but primarily designed for rendering individual objects or scenes with boundaries, posing challenges in handling unbounded scenes like the examples in Figure 5. To highlight the effectiveness of the proposed method, we particularly compare Vosh with VMesh on NeRF-Synthetic dataset [1]. The speed test in NeRF-Synthetic is based on the Lenovo Yoga 14s 2021 equipped with Iris Xe Graphics 80EU at a resolution of 800$\times$600. Due to the particularity of the NeRF-Synthetic dataset, the mesh extracted based on volume representation can usually achieve higher rendering quality than the volume representation itself after optimization [14]. So we mesh the density field of the entire voxel grid, which means that the $p_{remove}$ is fixed to $1$. Our method has better performance than VMesh in the aspect of rendering quality or speed, as shown in Table II for quantitative comparison and Figure 7 for qualitative comparison.
TABLE II: Quantitative evaluation of the proposed method and VMesh [20] in NeRF-Synthetic dataset [1].
Methods PSNR SSIM LPIPS FPS
VMesh 30.75 0.948 0.059 261.0
Vosh-Base 30.96 0.948 0.064 85.0
Vosh-Light 30.77 0.947 0.065 281.8
Figure 7: The rendering results and zoomed-in images on the NeRF-Synthetic dataset [1] from our method and VMesh [20].
Figure 8: The hybrid, mesh based and volume based rendering results obtained by our method and VMesh [20] on the NeRF-Synthetic dataset [1]. The PSNR in the figure represents the evaluation result of the current images.
We also provide the rendering results for the mesh part and the volume part separately, in addition to the rendering results of the hybrid representation of VMesh and Vosh, in Figure 8. Our reconstruction method of hybrid representation based on volumetric density field can restore more comprehensive scene information, thereby utilizing the extracted mesh to represent thin structures that are difficult to reconstruct in SDF, such as fishing nets and ropes in the figure. Hence, our method usually gains better rendering quality than VMesh.
VI-E Storage comparison
We conduct a statistical analysis of the graphics memory consumption and disk usage of the SOTA methods and the proposed method on the outdoor scenes of Mip-NeRF 360 [38], as shown in Table III. We use the occupying grid based approach proposed by SNeRG [8] to store voxels, and use methods similar to MobileNeRF [13] and NeRF2Mesh [14] to store mesh assets. Since we do not use the triplane representation which is the core contribution of MERF [18], but only recover the scene information and generate a hybrid representation based on the classic voxel representation SNeRG [8], it appears to be slightly larger than MERF in terms of disk storage. Although our method does not make any additional optimizations in terms of memory, it still has certain advantages compared to methods with similar rendering quality or speed.
TABLE III: Memory consumption and disk usage of the SOTA methods in outdoor scenes on the Mip-NeRF 360 dataset.
Methods VRAM (MB) ↓ Disk (MB) ↓
SNeRG++ 340 624
MERF 360 149
NeRF2Mesh 426 123
MobileNeRF 927 345
BakedSDF 953 807
3DGS 3701 1270
Vosh-Light 490 232
Vosh-Base 506 488
VI-F Ablation study
@@
We conduct ablation experiments on the meshing density field and voxel pruning parts of the hybrid representation on outdoor scenes of Mip-NeRF dataset [38], as shown in Table IV and Table V. The assessment involves a quantitative comparison, considering both rendering quality and speed.
We compare the variants of different optimization methods for differentiable surface optimization in Table IV. In addition to rendering quality, we also report the number of faces (in millions) of different variants . It can be seen that optimizing the geometry of the mesh based on differentiable rendering can significantly improve rendering quality. The experiments on surface operations indicate that it can better preserve rendering quality without mesh simplification, but might result in a large number of triangles. It might also cause some loss of rendering quality without performing mesh subdivision. So we use mesh subdivision and simplification based on rendering quality and normal angle change to significantly improve rendering quality while decreasing the number of triangles.
Compared with the mesh optimization in NeRF2Mesh [14], our method additionally considers the angle of normals between adjacent triangles during the optimization, which can improve rendering quality while reducing the number of triangles. At the same time, our meshing strategy based on rendering quality at the same stage also helps generating the high quality hybrid representation as shown in Fig 9.
TABLE IV: The ablation study in the differentiable surface refinement.
Variants PSNR SSIM LPIPS # faces
Mesh Baseline 22.04 0.475 0.448 1.76
Refine w/o subd 23.52 0.597 0.379 1.44
Refine w/o deci 23.61 0.601 0.377 2.14
Refine w/o norm 23.54 0.594 0.381 1.83
Refine 23.57 0.598 0.379 1.69
TABLE V: The ablation study in the construction of hybrid representation.
Variants PSNR SSIM LPIPS # faces # voxels
Meshing $p_{remove}=0.5$ 24.49 0.642 0.357 0.81 98999
$p_{remove}=1.0$ 23.95 0.600 0.384 1.69 47631
Refined mesh 23.57 0.598 0.379 1.69 0
Voxel grid 24.48 0.646 0.358 0 126533
Voxel prune. $\lambda_{voxel}=0$ 24.23 0.625 0.358 1.41 74317
$\lambda_{voxel}=0.01$ 23.90 0.606 0.377 1.41 36255
$r_{mesh}=256$ 23.80 0.604 0.377 1.41 41743
$r_{mesh}=512$ 23.94 0.609 0.373 1.41 52280
Vosh-Base 24.18 0.623 0.364 1.41 57438
Vosh-Light 23.74 0.602 0.377 1.41 2489
Figure 9: The hybrid, mesh based and volume based rendering results obtained by our method on Mip-NeRF 360 dataset [38]. The PSNR in the figure represents the evaluation result of the current images.
To better illustrate the effectiveness of each component in the construction process of our hybrid representation, we compare Vosh-Base with its variants in Table V, which includes the threshold $p_{remove}$ in meshing based on rendering quality, the weight of the voxel pruning loss $\lambda_{voxel}$, and the resolution $r_{mesh}$ for mesh occupancy grid. We also report the number of faces (in millions) that make up the mesh component and the number of alive voxels that make up the voxels component in the Vosh representation.
In the section of meshing density field, we compare the variants which $p_{remove}=0.5$ and $1.0$ with the Vosh-Base model with $p_{remove}=0.9$. Meanwhile, in addition to the rendering results of the initial voxel grid, we also provide refined rendering results for all surfaces obtained by meshing the entire voxel grid. When setting $p_{remove}=1.0$, it means that the entire density field is meshed, not based on rendering quality. The results show that this can only achieve a small improvement in quality. When setting $p_{remove}=0.5$, our hybrid representation that combines the advantages of voxel and mesh components achieve higher or comparable quality rendering results than the initial voxel grid.
@@
In the section of voxel pruning, we compare the variants $\lambda_{voxel}=0$ and $0.01$ with the Vosh-Base model with $\lambda_{voxel}=0.001$. We also list the rendering results of the variants $r_{mesh}=256$ and $512$, while Vosh-Base model turns off the check of mesh occupancy grid. Experimental results suggest that increasing the weight of voxel pruning loss or reducing the resolution of the mesh occupancy grid aids in compressing the number of alive voxels, leading to speed benefits at the expense of rendering quality.
At the same time, we also provide the hybrid, mesh based and volume based rendering results of several representative variants of Vosh, as shown in Figure 9. We use Vosh-Base in the second line as the benchmark, which sets $p_{remove}$ to 0.9 and $\lambda_{voxel}$ to 0.001. As a comparison, the first line only changes $p_{remove}$ to 0.5 and the third line only changes $\lambda_{voxel}$ to 0.01. On the basis of Vosh-Base, Vosh-Light sets $\lambda_{voxel}$ to 0.01, $p_{remove}$ to 1.0 and $r_{mesh}$ to 128. Based on the qualitative comparison in this figure, it can be seen that we can extract surfaces with higher reconstruction quality by determining the region where the density field meshes based on the difference in rendering quality through the threshold $p_{remove}$. And the balance between quality and speed in the hybrid representation model can be adjusted by setting the different weight $\lambda_{voxel}$ of the voxel pruning loss and the resolution $r_{mesh}$ for mesh occupancy grid.
VI-G Limitations